• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name sarty

    Log entry time 20:45:58 on July24,2006

    Entry number 175152


    keyword=A little trouble getting Q2 current set properly (HRS-L @ 343.64 GeV/c)

    Note: HRS-L is set as best as we can seem to get it for the needed "low momentum values" as per John Lerose's prescription for setting D, Q2, Q3 at 343.64 GeV/c. We find we can't get Q2 current to move from its current value of 148.0 A, but its field is only 0.4% away from the prescribed value. Yet, I am a bit worried that Q2 wasn't responding to setting in different current values, so I'm going to try completely resetting the HRS-L momentum by first putting the momentum value into the "P0 Set", seeing if all magnets start to respond (mostly Q2!)...if they do, I'll try to grab control of Q2 to get its field set properly.

    FIGURE 1